### v2

# Relatives of the Hermitian curve

Abstract
We introduce the notion of a relative of the Hermitian curve of degree $\sqrt{q}+1$ over $\mathbb{F}_{q}$, which is a plane curve defined by
$(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A{}^{t}\!(x,y,z)=0$
with $A\in GL(3,\mathbb{F}_{q})$, and study their basic properties together with various examples. One of the basic properties is that the number of $\mathbb{F}_{q}$-points of any relative of the Hermitian curve of degree $\sqrt{q}+1$ is congruent to $1$ modulo $\sqrt{q}$.
Key Words: Plane curve, Finite field, Rational point
MSC: 14G15, 14H50, 14G05, 11G20
1 Our object
Throughout this paper, $q$ denotes an even power of a prime number, so $\sqrt{q}$ is also a power of the prime number.
A Hermitian curve is a plane curve defined by
$(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A\begin{pmatrix}x\\ y\\ z\end{pmatrix}=0,$ (1)
where $A\in GL(3,\mathbb{F}_{q})$ satisfying the Hermitian condition ${}^{t}\!A^{(\sqrt{q})}=A$. Here the symbol $A^{(\sqrt{q})}$ means taking $\sqrt{q}$-th power on each entry of $A$.
Forgetting the Hermitian condition, the equation (1) still defines a nonsingular plane curves of degree $\sqrt{q}+1$ over $\mathbb{F}_{q}$, which will be denoted by $C_{A}$. We call such a curve a relative of the Hermitian curve, or shortly, a Hermitian-relative curve, and want to study them.
All Hermitian-relative curves, including Hermitian curves themselves, are projectively equivalent each other over the algebraic closure $\overline{\mathbb{F}}_{q}$ of $\mathbb{F}_{q}$, which is due to Pardini [7, Proof of Proposition 3.7]. We will give a proof of this fact in the next section for the convenience of readers.
Any Hermitian-relative curve $C$ over $\mathbb{F}_{q}$ has a particular property on the number $N_{q}(C)$ of $\mathbb{F}_{q}$-points:
$N_{q}(C)\equiv 1\mod\sqrt{q}.$
The proof of this fact is also given in the following section.
Since the equation (1) is homogeneous, the curve $C_{A}$ depends only on the image $\overline{A}$ of $A$ under the canonical map $GL(3,\mathbb{F}_{q})\to PGL(3,\mathbb{F}_{q})$. By abuse of notation, $\overline{A}$ will be frequently denoted by $A$, namely an element of $GL(3,\mathbb{F}_{q})$ and the image of the element in $PGL(3,\mathbb{F}_{q})$ will be used interchangeably. So the equation $A=B$ in $PGL(3,\mathbb{F}_{q})$ means that there exists an element $\lambda\in\mathbb{F}_{q}^{\ast}$ so that $A=\lambda B$ in $GL(3,\mathbb{F}_{q})$.
Remark 1.1
If $\overline{A}\in PGL(3,\mathbb{F}_{q})$ satisfies the Hermitian condition ${}^{t}\overline{A}^{(\sqrt{q})}=\overline{A}$, then we can choose an element $A\in GL(3,\mathbb{F}_{q})$ satisfying the Hermitian condition whose image in $PGL(3,\mathbb{F}_{q})$ is the assigned $\overline{A}$.
Proof.
Let $A$ be any source of $\overline{A}$. Then there is an element $\lambda\in\mathbb{F}_{q}^{\ast}$ such that ${}^{t}A^{(\sqrt{q})}=\lambda A$. Taking $\sqrt{q}$-th power and transpose of the both side, $A=\lambda^{\sqrt{q}}\left({}^{t}\!A^{(\sqrt{q})}\right)$. So $A=\lambda^{\sqrt{q}+1}A$, which implies $\lambda^{\sqrt{q}+1}=1$. Therefore there is an element $\rho\in\mathbb{F}_{q}$ so that $\rho^{1-\sqrt{q}}=\lambda$ (which is a special case of Hilbert’s Theorem 90). Then $\overline{\rho A}=\overline{A}$ and $\rho A$ satisfies the Hermitian condition. ∎
Annotation
Although a homogeneous coordinates of a point has an ambiguity by nonzero multiple, hereafter, for any $\mathbb{F}_{q}$-point, we understand all values of the coordinates are elements of $\mathbb{F}_{q}$.
2 Basic properties
The polynomial of type (1) is a linear combination of the monomials
$x^{\sqrt{q}+1},y^{\sqrt{q}+1},z^{\sqrt{q}+1},x^{\sqrt{q}}y,y^{\sqrt{q}}z,z^{\sqrt{q}}x,xy^{\sqrt{q}},yz^{\sqrt{q}},zx^{\sqrt{q}}$
over $\mathbb{F}_{q}$, and vice versa. So those polynomials forms an $\mathbb{F}_{q}$-vector space of dimension $9$. Moreover the vector space is stable under the linear transformations on coordinates $x,y,z$. We should note that by a linear transformation
$\begin{pmatrix}x\\ y\\ z\end{pmatrix}\mapsto T\begin{pmatrix}x\\ y\\ z\end{pmatrix}$
with $T\in PGL(3,\mathbb{F}_{q})$, the polynomial $(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A\begin{pmatrix}x\\ y\\ z\end{pmatrix}$ goes to
$(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}}){}^{t}T^{(\sqrt{q})}AT\begin{pmatrix}x\\ y\\ z\end{pmatrix}.$
The properties mentioned above hold true after any field extension $\mathbb{F}_{q}$.
For given two matrices $A,B\in GL(3,\mathbb{F}_{q})$ (or in $PGL(3,\mathbb{F}_{q})$), we say that $A$ and $B$ are equivalent if two curves $C_{A}$ and $C_{B}$ are projectively equivalent over $\mathbb{F}_{q}$. This equivalence relation is denoted by $A\sim B$. In other words, $A\sim B$ if there is a matrix $T\in GL(3,\mathbb{F}_{q})$ and $\lambda\in\mathbb{F}_{q}^{\ast}$ such that $\lambda B={}^{t}T^{(\sqrt{q})}AT$ (or $B={}^{t}T^{(\sqrt{q})}AT$ in $PGL(3,\mathbb{F}_{q})$).
Remark 2.1
Even if $\det A=0$, the curve defined by (1) makes sense; however the curve is nonsingular if and only if $\det A\neq 0$. Actually, the partial derivatives of $f=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A\begin{pmatrix}x\\ y\\ z\end{pmatrix}$ are given by
$(f_{x},f_{y},f_{z})=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A.$ (2)
Proposition 2.2
For $A\in GL(3,\mathbb{F}_{q})$, the dual curve of $C_{A}$ is projectively equivalent to $C_{A^{(\sqrt{q})}}$ over $\mathbb{F}_{q}$.
Proof.
Let $u,v,w$ be the dual coordinate of the dual projective plane with point-line incidence $ux+vy+wz=0$. By (2), the equation after eliminating $x,y,z$ from
$(u,v,w)=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A$ (3)
and (1) is an equation of the dual curve. Taking entry-wise $\sqrt{q}$-th power of (1),
$(x^{q},y^{q},z^{q})A^{(\sqrt{q})}\begin{pmatrix}x^{\sqrt{q}}\\ y^{\sqrt{q}}\\ z^{\sqrt{q}}\end{pmatrix}=0,$
which can be written
$(u^{\sqrt{q}},v^{\sqrt{q}},w^{\sqrt{q}}){}^{t}\!A^{-1}\begin{pmatrix}u\\ v\\ w\end{pmatrix}=0$
by (3). On the other hand, since
(${}^{t}\!\left({}^{t}\!A\right)^{(\sqrt{q})}\left({}^{t}\!A^{-1}\right){}^{t}\!A=A^{(\sqrt{q})},$
${}^{t}\!A^{-1}\sim A^{(\sqrt{q})}$. This completes the proof. ∎
Next is the first main purpose of this section.
Theorem 2.3
Let $f(x,y,z)\in\mathbb{F}_{q}[x,y,z]$ be a homogeneous polynomial of degree $\sqrt{q}+1$. The curve $C$ defined by $f(x,y,z)=0$ in $\mathbb{P}^{2}$ is projectively equivalent over $\overline{\mathbb{F}}_{q}$ to a Hermitian curve of degree $\sqrt{q}+1$ if and only if $C$ is a Hermitian-relative curve, that is, there exists $A\in PGL(3,\mathbb{F}_{q})$ such that
$f(x,y,z)=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A\begin{pmatrix}x\\ y\\ z\end{pmatrix}.$
Proof.
(Pardini) First suppose that $C$ is projectively equivalent to
$g(x,y,z)=x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+z^{\sqrt{q}+1}=0$
over $\overline{\mathbb{F}}_{q}$. Then there is a matrix $T=(t_{i,j})\in GL(3,\overline{\mathbb{F}}_{q})$ so that
$f(x,y,z)=g((x,y,z){}^{t}T),$
that is,
$f(x,y,z)$ $=\sum_{j=1}^{3}\left(t_{j1}x+t_{j2}y+t_{j3}z\right)^{\sqrt{q}+1}$
$=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}}){}^{t}T^{(\sqrt{q})}T\begin{pmatrix}x\\ y\\ z\end{pmatrix}.$
Note that each entry of ${}^{t}T^{(\sqrt{q})}T$ gives a coefficient of distinct monomial of $f(x,y,z)$. So ${}^{t}T^{(\sqrt{q})}T\in PGL(3,\mathbb{F}_{q})$.
To see the converse, let us consider a morphism associated with an assigned $A\in PGL(3,\mathbb{F}_{q})$:
$\Phi_{A}:PGL(3,\overline{\mathbb{F}}_{q})\ni T\mapsto{}^{t}T^{(\sqrt{q})}AT\in PGL(3,\overline{\mathbb{F}}_{q}).$
If $\Phi_{A}(S)=\Phi_{A}(T)$ for $S,T\in PGL(3,\overline{\mathbb{F}}_{q})$, then
(${}^{t}(ST^{-1})^{(\sqrt{q})}A(ST^{-1})=A$
which means $ST^{-1}$ gives a linear automorphism of $C_{A}$ over $\overline{\mathbb{F}}_{q}$ in $\mathbb{P}^{2}$.
The automorphism group of $C_{A}$ over $\overline{\mathbb{F}}_{q}$ is finite if $\sqrt{q}>2$, because the genus of $C_{A}$ is grater than $1$. When $\sqrt{q}=2$, $C_{A}$ is an elliptic curve over $\overline{\mathbb{F}}_{q}$ in $\mathbb{P}^{2}$. The linear automorphism group of any elliptic curve in $\mathbb{P}^{2}$ is also finite. Therefore the image of $\Phi_{A}$ is a constructible set (which is a theorem of Chevalley [6, Corollary 2 in I §8]) of dimension $8$ in the target space $PGL(3,\overline{\mathbb{F}}_{q})$. Therefore two images of $\Phi_{A}$ and $\Phi_{B}$ meet in $PGL(3,\overline{\mathbb{F}}_{q})$ for any two $A,B\in PGL(3,\mathbb{F}_{q})$, that is, there is an element $T\in PGL(3,\overline{\mathbb{F}}_{q})$ such that ${}^{t}T^{(\sqrt{q})}AT=B$. In particular, if one takes as $B=I_{3}$, one gets a projective transformation over $\overline{\mathbb{F}}_{q}$ which sends $C_{A}$ to a Hermitian curve. ∎
In the next corollary, $T_{P}(C)$ denotes the (embedded) tangent line to a plane curve $C$ at $P$, and $i(C.T_{P}(C);P)$ the intersection multiplicity of $C$ and $T_{P}(C)$ at $P$.
Corollary 2.4
Suppose that $q>4$. Among nonsingular plane curves of degree $\sqrt{q}+1$, a Hermitian-relative curve $C$ is characterized by the property: for any $\overline{\mathbb{F}}_{q}$-point $P$ of $C$, $i(C.T_{P}(C);P)$ is either $\sqrt{q}$ or $\sqrt{q}+1$.
Proof.
This characterization comes from [4, Theorem 6.1, and Theorem 3.5] together with Theorem 2.3
The following is the second main purpose of this section.
Theorem 2.5
Let $C$ be a relative of the Hermitian curve of degree $\sqrt{q}+1$ over $\mathbb{F}_{q}$. Then
$N_{q}(C)\equiv 1\mod\sqrt{q}.$
In particular, the set $C(\mathbb{F}_{q})$ of $\mathbb{F}_{q}$-points on $C$ is not empty.
Proof.
Let $g=\frac{1}{2}\sqrt{q}(\sqrt{q}-1)$, which is the genus of any Hermitian-relative curve. Let $C^{\prime}$ be a Hermitian curve of degree $\sqrt{q}+1$, e.g., $C^{\prime}$ is one defined by $x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+z^{\sqrt{q}+1}=0.$ By Theorem 2.3, $C$ and $C^{\prime}$ are projectively equivalent over some extension field $\mathbb{F}_{q^{s}}$. Let
$Z_{C}(t)=\prod_{j=1}^{2g}(1-\alpha_{j}t)/(1-t)(1-qt)$
be the zeta function of $C$ over $\mathbb{F}_{q}$. By Weil’s theorem,
$N_{q^{s}}(C)=1+q^{s}-\sum_{j=1}^{2g}\alpha_{j}^{s}\text{\ and \ }\alpha_{j}=\sqrt{q}\eta_{j}\text{\ with \ }|\eta_{j}|=1.$ (4)
On the other hand, since $C^{\prime}$ is a maximal curve over $\mathbb{F}_{q}$, its zeta function is
$Z_{C^{\prime}}(t)=\left(1-(-\sqrt{q})t\right)^{2g}/(1-t)(1-qt).$
Since $C$ and $C^{\prime}$ are isomorphic over $\mathbb{F}_{q^{s}}$,
$\{\alpha_{1}^{s},\dots,\alpha_{2g}^{s}\}=\{\overbrace{(-\sqrt{q})^{s},\dots,(-\sqrt{q})^{s}}^{2g}\}.$
(As for this classical theory, one may consult [9, Chapter 5].)
Since we may choose $s$ to be even, $\eta_{j}^{s}=1$ for each $j=1,\dots,2g$. In particular, $\eta_{j}$ is an algebraic integer, and so is $\sum_{j=1}^{2g}\eta_{j}$. Going back to the formula (4) (with $s=1$),
$\sum_{j=1}^{2g}\eta_{j}=\frac{q+1-N_{q}(C)}{\sqrt{q}}.$
Since $\sqrt{q}$ is an integer by our assumption, $\sum_{j=1}^{2g}\eta_{j}$ is a rational number. Therefore it is a rational integer. Hence
$N_{q}(C)=q+1-\sqrt{q}(\sum_{j=1}^{2g}\eta_{j})\equiv 1\mod\sqrt{q}.$
This completes the proof. ∎
Corollary 2.6
For a relative of the Hermitian curve $C$ over $\mathbb{F}_{q}$, there is a nonnegative integer $m=m_{C}$ with
$m=q\text{\ or \ }\sqrt{q}+2\geq m\geq 0$
so that $N_{q}(C)=m\sqrt{q}+1$. Moreover $m_{C}=q$ if and only if either $C$ is Hermitian, or $q=4$ and $C$ is projectively equivalent to the curve
$x^{\sqrt{q}+1}+\omega y^{\sqrt{q}+1}+\omega^{2}z^{\sqrt{q}+1}=0.$
over $\mathbb{F}_{4}$, where $\omega\in\mathbb{F}_{4}\setminus\mathbb{F}_{2}$.
Proof.
The additional statement is somewhat classical [3, 8, 5]. For the first statement, when $q=4$, there is nothing new from Theorem 2.5. Hence we assume that $q>4$. Let $C=C_{A}$, and $(x_{0},y_{0},z_{0})\in C(\mathbb{F}_{q})$. Then $(x_{0}^{\sqrt{q}},y_{0}^{\sqrt{q}},z_{0}^{\sqrt{q}})A\begin{pmatrix}x_{0}\\ y_{0}\\ z_{0}\end{pmatrix}=0.$ Taking $\sqrt{q}$-th power, and noting $x_{0}^{q}=x_{0},y_{0}^{q}=y_{0},z_{0}^{q}=z_{0}$, we know that $(x_{0},y_{0},z_{0})\in C_{{}^{t}\!A^{(\sqrt{q})}}$. If $C$ is not Hermitian, $C_{A}$ and $C_{{}^{t}\!A^{(\sqrt{q})}}$ are distinct irreducible curves. Hence the intersection of these two curves consists of at most $(\sqrt{q}+1)^{2}$ points, which gives an upper bound for $N_{q}(C)$. ∎
3 Example I - diagonal case
Remaining three sections, selected examples are explained; however, the division of sections is rather superficial (cf. Example 3.4 and Example 4.1). In order to compute examples, some notes will be explained. The trace $\mathbb{F}_{q}\ni x\mapsto x^{\sqrt{q}}+x\in\mathbb{F}_{\sqrt{q}}$ and the norm $\mathbb{F}_{q}^{\ast}\ni x\mapsto x^{\sqrt{q}+1}\in\mathbb{F}_{\sqrt{q}}^{\ast}$ are indicated by $\operatorname{Tr}$ and $\operatorname{Nm}$ respectively.
Lemma 3.1
There are two exact sequences:
$\begin{array}[]{rcccccccl}0\to&\mathbb{F}_{\sqrt{q}}&\stackrel{{\scriptstyle i}}{{\to}}&\mathbb{F}_{q}&\to&\mathbb{F}_{q}&\xrightarrow{\operatorname{Tr}}&\mathbb{F}_{\sqrt{q}}&\to 0\\ &&&\alpha&\mapsto&\alpha^{\sqrt{q}}-\alpha&&&\end{array};$ (5)
$\begin{array}[]{rcccccccl}1\to&\mathbb{F}_{\sqrt{q}}^{\ast}&\stackrel{{\scriptstyle i}}{{\to}}&\mathbb{F}_{q}^{\ast}&\to&\mathbb{F}_{q}^{\ast}&\xrightarrow{\operatorname{Nm}}&\mathbb{F}_{\sqrt{q}}^{\ast}&\to 1\\ &&&\alpha&\mapsto&\alpha^{\sqrt{q}-1}&&&\end{array},$ (6)
where $i$ means natural inclusion.
Proof.
This lemma is a special case of Hilbert’s Theorem 90, and the proof is easy. ∎
Corollary 3.2
  1. (i)
    The equation of $X$ with $\beta\in\mathbb{F}_{q}$
    $X^{\sqrt{q}}-X-\beta=0$ (7)
[…]
  2. (ii)
    The equation of $X$ with $\beta\in\mathbb{F}_{q}^{\ast}$
    $X^{\sqrt{q}-1}-\beta=0$ (8)
[…]
Proof.
(i) Let $\alpha\in\mathbb{F}_{q}$ be a root of (7). Then $\beta=\alpha^{\sqrt{q}}-\alpha$. Hence $\operatorname{Tr}\beta=0$ by (5). Conversely, if $\operatorname{Tr}\beta=0$, there is an element $\alpha\in\mathbb{F}_{q}$ such that $\alpha^{\sqrt{q}}-\alpha=\beta$ by (5) again. When $\alpha$ is a root of (7), the set of all roots is $\{\alpha+\gamma|\gamma\in\mathbb{F}_{\sqrt{q}}\}$.
(ii) can be proved by using (6). ∎
In this section, we handle only diagonal matrices; $\operatorname{diag}[a,b,c]$ denotes the diagonal matrix whose diagonal entries are $a,b,c$ from the upper left to the lower right.
Let $A\in PGL(3,\mathbb{F}_{q})$, which is diagonal. Then $A$ is equivalent to one of the following three types.
Example 3.3
(Type 0) $A\sim\operatorname{diag}[1,1,1]$. So the corresponding curve $C_{A}$ is projectively equivalent to the Hermitian curve $x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+z^{\sqrt{q}+1}=0$, and $N_{q}(C_{A})=(\sqrt{q})^{3}+1$. Hence $m_{C_{A}}=q$.
Example 3.4
(Type 1) $A\sim\operatorname{diag}[1,1,\omega]$ for some $\omega\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$. Hence $C_{A}$ is projectively equivalent to the curve $D_{\omega}$ defined by
$x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+\omega z^{\sqrt{q}+1}=0.$ (9)
If $(x_{0},y_{0},z_{0})\in(\mathbb{F}_{q})^{3}$ is a solution of (9), then it is a solution of the simultaneous equations:
[…]
because $x_{0}^{\sqrt{q}+1},y_{0}^{\sqrt{q}+1},z_{0}^{\sqrt{q}+1}\in\mathbb{F}_{\sqrt{q}}$ and $\{1,\omega\}$ are linearly independent over $\mathbb{F}_{\sqrt{q}}$. Therefore
$D_{\omega}(\mathbb{F}_{q})=\{(\zeta,1,0)|\operatorname{Nm}\zeta=-1\}.$
Hence $N_{q}(C_{A})=\sqrt{q}+1$, and $m_{C_{A}}=1$.
Example 3.5
(Type 2) $A\sim\operatorname{diag}[1,\omega,\omega^{\prime}]$ for some $\omega,\omega^{\prime}\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$ with $\omega(\omega^{\prime})^{-1}\not\in\mathbb{F}_{\sqrt{q}}^{\ast}$. Hence $C_{A}$ is projectively equivalent the curve $D_{\omega,\omega^{\prime}}$ defined by
$x^{\sqrt{q}+1}+\omega y^{\sqrt{q}+1}+\omega^{\prime}z^{\sqrt{q}+1}=0.$ (10)
Since any two elements of $\{1,\omega,\omega^{\prime}\}$ are linearly independent over $\mathbb{F}_{\sqrt{q}}$, there is a unique relation up to $\mathbb{F}_{\sqrt{q}}^{\ast}$-multiple among those three elements, say
$\alpha+\beta\omega+\gamma\omega^{\prime}=0.$
Since none of three coefficients $\alpha,\beta,\gamma$ are 00, we may assume that $\alpha=1$, and the first coordinate of any solution of (10) is also $1$. Then $D_{\omega,\omega^{\prime}}(\mathbb{F}_{q})=\{(1,\zeta,\eta)|\operatorname{Nm}\zeta=\beta,\operatorname{Nm}\eta=\gamma\}$. Hence $N_{q}(C_{A})=(\sqrt{q}+1)^{2}$, and hence $m_{C}=\sqrt{q}+2$.
4 Example II - non-diagonal case
The following examples, including the diagonal case explained above, are inspired by Hirschfeld’s classification of cubic plane curves over any finite fields [2, Section 11]. Actually, the family of supersingular plane cubic curves over $\mathbb{F}_{4}$ coincides with the family of relatives of the Hermitian curve over $\mathbb{F}_{4}$, because any curves in each family is characterized by disappearance of the term $xyz$ in its equation (see, [1, IV, Prop. 4.21]).
Example 4.1
Let $\omega\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$, and
$A=\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&\omega\end{pmatrix}.$
Then $C_{A}$ is defined by
$x^{\sqrt{q}}y+xy^{\sqrt{q}}+\omega z^{\sqrt{q}+1}=0.$
The curve $C_{A}$ is projectively equivalent over $\mathbb{F}_{q}$ to a diagonal curve of type 1. In fact, choose $u\in\mathbb{F}_{q}$ so that $\operatorname{Tr}u=1$ and $\alpha\in\mathbb{F}_{q}^{\ast}$ so that $\operatorname{Nm}\alpha=-1$. Let $T_{0}=\begin{pmatrix}u&\alpha u^{\sqrt{q}}&0\\ 1&-\alpha&0\\ 0&0&1\end{pmatrix}.$ Then it is easy to see that
${}^{t}\!T_{0}^{(\sqrt{q})}\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&\omega\end{pmatrix}T_{0}=\operatorname{diag}[1,1,\omega].$
Example 4.2
Let us consider the matrix $A=\begin{pmatrix}1&0&0\\ 0&\omega&1\\ 0&1&0\end{pmatrix},$ where $\omega\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$. So the curve $C_{A}$ is defined by
$x^{\sqrt{q}+1}+\omega y^{\sqrt{q}+1}+y^{\sqrt{q}}z+yz^{\sqrt{q}}=0.$ (11)
Proposition 4.3
For the curve (11), $C_{A}(\mathbb{F}_{q})=\{(0,0,1)\}$, and hence $m_{C_{A}}=0$.
Proof.
For any $(x_{0},y_{0},z_{0})\in C_{A}(\mathbb{F}_{q})$, three elements $x_{0}^{\sqrt{q}+1}$, $y_{0}^{\sqrt{q}+1}$, and $y_{0}^{\sqrt{q}}z_{0}+y_{0}z_{0}^{\sqrt{q}}=\operatorname{Tr}(y_{0}^{\sqrt{q}}z_{0})$ are in $\mathbb{F}_{\sqrt{q}}$. Since $\{\omega,1\}$ are linearly independent over $\mathbb{F}_{\sqrt{q}}$, we have
$x_{0}^{\sqrt{q}+1}+y_{0}^{\sqrt{q}}z_{0}+y_{0}z_{0}^{\sqrt{q}}=0\textrm{\ and \ }y_{0}^{\sqrt{q}+1}=0.$
Hence $x_{0}=y_{0}=0$. ∎
Example 4.4
Next we consider the matrix $A=\begin{pmatrix}\omega&0&0\\ 0&\omega&1\\ 0&1&0\end{pmatrix},$ where $\omega\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$. The corresponding curve $C_{A}$ is defined by
$\omega x^{\sqrt{q}+1}+\omega y^{\sqrt{q}+1}+y^{\sqrt{q}}z+yz^{\sqrt{q}}=0.$ (12)
Proposition 4.5
For the curve (12),
$\begin{split}C_{A}(\mathbb{F}_{q})=\{P_{\zeta}=(\zeta,1,0)|&\zeta^{\sqrt{q}+1}=-1\}\cup\{R=(0,0,1)\}\\ &\cup\{Q_{\xi\rho}=(\xi,\rho,1)|\rho^{\sqrt{q}-1}=-1,\ \xi^{\sqrt{q}+1}=\rho^{2}\}.\end{split}$
In particular, $N_{q}(C_{A})=q+\sqrt{q}+1$ and $m_{C_{A}}=\sqrt{q}+1$.
Proof.
On the line $\{z=0\}$ [^1] During computation of examples, frequently use this type of abbreviation., the equation is $\omega(x^{\sqrt{q}+1}+y^{\sqrt{q}+1})=0.$ Hence
$C_{A}(\mathbf{F}_{q})\cap\{z=0\}=\{P_{\zeta}|\zeta^{\sqrt{q}+1}=-1.\}$
In affine space $\mathbb{P}^{2}\setminus\{z=0\}$, we may assume that the $z$-coordinate is always $1$. If $(x_{0},y_{0},1)$ is an $\mathbb{F}_{q}$-point of $C_{A}$, it satisfies
$\omega(x_{0}^{\sqrt{q}+1}+y_{0}^{\sqrt{q}+1})+y_{0}^{\sqrt{q}}+y_{0}=0,$ (13)
and vice versa. Since $x_{0}^{\sqrt{q}+1}+y_{0}^{\sqrt{q}+1}$ and $y_{0}^{\sqrt{q}}+y_{0}$ are elements of $\mathbb{F}_{\sqrt{q}}$, and $\{\omega,1\}$ are linearly independent over $\mathbb{F}_{\sqrt{q}}$, (13) is equivalent to
[…]
Hence
$C_{A}(\mathbb{F}_{q})\cap\{z\neq 0\}=\{R\}\cup\{Q_{\xi\rho}|\rho^{\sqrt{q}-1}=-1,\ \xi^{\sqrt{q}+1}=\rho^{2}\}.$
It is easy to show that |$|\{P_{\xi}\}|=\sqrt{q}+1$, and |$|\{R\}\cup\{Q_{\xi\rho}\}|=1+(\sqrt{q}-1)(\sqrt{q}+1).$
Example 4.6
Let $A=\begin{pmatrix}1&0&0\\ 1&0&1\\ 0&1&0\end{pmatrix},$ Hence the curve $C_{A}$ is defined by
$x^{\sqrt{q}+1}+xy^{\sqrt{q}}+y^{\sqrt{q}}z+yz^{\sqrt{q}}=0.$
Proposition 4.7
For the curve $C_{A}$ in Example 4.6, we have
$N_{q}(C_{A})=q+1,\text{\ and so\ }m_{C_{A}}=\sqrt{q}.$
Proof.
First consider the case $x=0$. Then the equation becomes $y^{\sqrt{q}}z+yz^{\sqrt{q}}=0$, which has obviously $\sqrt{q}+1$ homogeneous solutions that are $\mathbb{F}_{q}$points.
Next suppose that $x=1$. Then the equation becomes
$01+y^{\sqrt{q}}+y^{\sqrt{q}}z+yz^{\sqrt{q}}=$ 0 . (14)
Let $(y_{0},z_{0})$ be any $\mathbb{F}_{q}$-solution of (14). Since $y_{0}^{\sqrt{q}}z_{0}+y_{0}z_{0}^{\sqrt{q}}=\operatorname{Tr}(y_{0}^{\sqrt{q}}z_{0})\in\mathbb{F}_{\sqrt{q}}$, the remaining term $1+y_{0}^{\sqrt{q}}$ is also an element of $\mathbb{F}_{\sqrt{q}}$, and so is $y_{0}^{\sqrt{q}}$. Hence $y_{0}=y_{0}^{q}=(y_{0}^{\sqrt{q}})^{\sqrt{q}}\in\mathbb{F}_{\sqrt{q}}$. Therefore the point $(y_{0},z_{0})$ satisfies
$01+y_{0}+(z_{0}+z_{0}^{\sqrt{q}})y_{0}=$ 0 (15)
from (14). Put $z_{0}+z_{0}^{\sqrt{q}}=t_{0}$, which is an elements of $\mathbb{F}_{\sqrt{q}}$. If $t_{0}\neq-1$, $y_{0}=\frac{-1}{t_{0}+1}$, which is uniquely determined by $t_{0}$. For each $t_{0}\in\mathbb{F}_{\sqrt{q}}\setminus\{-1\}$, we can find $\sqrt{q}$ solutions of $z+z^{\sqrt{q}}=t_{0}$ in $\mathbb{F}_{q}$. If $t_{0}=-1$, There are no solutions of (15) on $y_{0}$. Hence $|C_{A}(\mathbb{F}_{q})\cap\{x\neq 0\}|=\sqrt{q}(\sqrt{q}-1).$ Therefore the number of $\mathbb{F}_{q}$-points of $C_{A}$ is totally
$(\sqrt{q}+1)+\sqrt{q}(\sqrt{q}-1)=q+1.$
This completes the proof. ∎
5 Example III - non-diagonal case (2)
In this section, we handle the curve
$x^{\sqrt{q}}z+xy^{\sqrt{q}}+\omega yz^{\sqrt{q}}=0,$ (16)
where $\omega\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$. The number of $\mathbb{F}_{q}$-points of curves in each examples explained in previous two sections does not depend on the choice of $\omega\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$, however that of the curve (16) depends on the choice of $\omega$. So we denote the corresponding matrix $\begin{pmatrix}0&0&1\\ 1&0&0\\ 0&\omega&0\end{pmatrix}$ by $A_{\omega}$.
Lemma 5.1
Let $C_{A_{\omega}}$ be a curve defined by (16).
(i)   $C_{A_{\omega}}(\mathbb{F}_{q})\cap\{z=0\}=\{(1,0,0),(0,1,0)\}.$
${\rm(ii)}\begin{split}(C_{A_{\omega}}(\mathbb{F}_{q})\cap&\{z\neq 0\})\cap\{y^{\sqrt{q}+1}\neq 1\}\\ &=\left\{\left(\frac{\omega^{\sqrt{q}}y_{0}^{\sqrt{q}}-\omega y_{0}^{2}}{y_{0}^{\sqrt{q}+1}-1},y_{0},1\right)\mid y_{0}\in\mathbb{F}_{q},\ y_{0}^{\sqrt{q}+1}-1\neq 0\right\}.\end{split}$
In particular,
$N_{q}(C_{A_{\omega}})\geq q-\sqrt{q}+1.$
Proof.
The assertion (i) is obvious. Let $(x_{0},y_{0},1)\in C_{A_{\omega}}(\mathbb{F}_{q})\cap\{z\neq 0\}$. Then
$x_{0}^{\sqrt{q}}+x_{0}y_{0}^{\sqrt{q}}+\omega y_{0}=0.$ (17)
Taking $\sqrt{q}$-th power of this equation, we have
$x_{0}+x_{0}^{\sqrt{q}}y_{0}+\omega^{\sqrt{q}}y_{0}^{\sqrt{q}}=0.$ (18)
[…]
$x_{0}(y_{0}^{\sqrt{q}+1}-1)-(\omega^{\sqrt{q}}y_{0}^{\sqrt{q}}-\omega y_{0}^{2})=0.$
Hence, when $y_{0}^{\sqrt{q}+1}-1\neq 0$, $x_{0}=\frac{\omega^{\sqrt{q}}y_{0}^{\sqrt{q}}-\omega y_{0}^{2}}{y_{0}^{\sqrt{q}+1}-1}$. Conversely, by straightforward computation, we get
$(\frac{\omega^{\sqrt{q}}y_{0}^{\sqrt{q}}-\omega y_{0}^{2}}{y_{0}^{\sqrt{q}+1}-1})^{\sqrt{q}}+(\frac{\omega^{\sqrt{q}}y_{0}^{\sqrt{q}}-\omega y_{0}^{2}}{y_{0}^{\sqrt{q}+1}-1})y_{0}^{\sqrt{q}}+\omega y_{0}=0.$
Putting (i), (ii) together, we have $q-(\sqrt{q}+1)+2$ points over $\mathbb{F}_{q}$. ∎
For computing the number of $\mathbb{F}_{q}$-points of $(C_{A_{\omega}}(\mathbb{F}_{q})\cap\{z\neq 0\})\cap\{y^{\sqrt{q}+1}=0\},$ we fix a primitive element $\omega_{0}\in\mathbb{F}_{q}^{\ast}$.
Note that $\omega_{0}^{s}\in\mathbb{F}_{\sqrt{q}}^{\ast}$ if and only if $s\equiv 0\mod\sqrt{q}+1$.
Lemma 5.2
(1) Suppose that $q$ is odd. Then
(i) $-1=\omega_{0}^{\frac{q-1}{2}},$ and
(ii) $\ker\operatorname{Tr}=\{0\}\cup\{\omega_{0}^{\frac{\sqrt{q}+1}{2}+s(\sqrt{q}+1)}\mid s=0,1,\dots,\sqrt{q}-2\}.$
(2) Suppose that $q$ is even. Then
(iii) $\ker\operatorname{Tr}=\{0\}\cup\{\omega_{0}^{s(\sqrt{q}+1)}\mid s=0,1,\dots,\sqrt{q}-2\}$.
Proof.
(i) Since $\omega_{0}$ is of order $q-1$ in the multiplicative group $\mathbb{F}_{q}^{\ast}$ and $q-1$ is even, $\omega_{0}^{\frac{q-1}{2}}=-1$.
(ii) For $b\in\mathbb{F}_{q}^{\ast}$, $\operatorname{Tr}b=0$ means $b^{\sqrt{q}-1}=-1$. Letting $b=\omega_{0}^{l}$, we have $\omega_{0}^{l(\sqrt{q}-1)}=\omega_{0}^{\frac{q-1}{2}}$. Hence $l(\sqrt{q}-1)\equiv\frac{\sqrt{q}+1}{2}(\sqrt{q}-1)\mod q-1$, which means that $l\equiv\frac{\sqrt{q}+1}{2}\mod\sqrt{q}+1$.
(iii) Similar to the case (ii). ∎
Theorem 5.3
  1. (1)
    If $(\sqrt{q}+1,3)=1$, then
    $N_{q}(C_{A_{\omega}})=q+1\text{\ and so \ }m_{C_{A_{\omega}}}=\sqrt{q}.$
  2. (2)
    When $\sqrt{q}+1$ is divisible by $3$,
    1. (a)
      if $\omega^{\frac{\sqrt{q}+1}{3}}\in\mathbb{F}_{\sqrt{q}}^{\ast}$, then
      $N_{q}(C_{A_{\omega}})=q+2\sqrt{q}+1\text{\ and so \ }m_{C_{A_{\omega}}}=\sqrt{q}+2,\text{\ and}$
    2. (b)
      if $\omega^{\frac{\sqrt{q}+1}{3}}\not\in\mathbb{F}_{\sqrt{q}}^{\ast}$, then
      $N_{q}(C_{A_{\omega}})=q-\sqrt{q}+1\text{\ and \ }m_{C_{A_{\omega}}}=\sqrt{q}-1.$
Proof.
We have to count the number of $\mathbb{F}_{q}$-solutions of the simultaneous equations
[…]
By changing the sign $y\mapsto-y,$ this system is equivalent to
[…]
For a given $y_{0}\in\mathbb{F}_{q}^{\ast}$ with $y_{0}^{\sqrt{q}+1}=1$, we ask whether the equation on $x$
$x^{\sqrt{q}}-y_{0}^{\sqrt{q}}x-\omega y_{0}=0$ (19)
has an $\mathbb{F}_{q}$-solution or not. Since $\operatorname{Nm}y_{0}^{\sqrt{q}}=\operatorname{Nm}y_{0}=1$, there is an element $c\in\mathbb{F}_{q}^{\ast}$ such that $c^{1-\sqrt{q}}=y_{0}^{\sqrt{q}}$ by Lemma 3.1. Therefore the problem above is transformed whether the equation on $cx$
$(cx)^{\sqrt{q}}-(cx)-\omega c^{\sqrt{q}}y_{0}=0$ (20)
has an $\mathbb{F}_{q}$-solution or not. From Corollary 3.2, (20) has an $\mathbb{F}_{q}$-solution if and only if $\operatorname{Tr}(\omega c^{\sqrt{q}}y_{0})=0$, and in this case all solutions of (20) are $\mathbb{F}_{q}$-solutions.
Let $i,j,k$ be integers with $11\leq i,j,k\leq q-$ 1 satisfying
$\omega=\omega_{0}^{k},\ c=\omega_{0}^{j},\text{\ and \ }y_{0}=\omega_{0}^{i}.$
Since $y_{0}^{\sqrt{q}+1}=1$, there is a unique integer $l$ with $11\leq l\leq\sqrt{q}+$ 1 so that $i\equiv l(\sqrt{q}-1)\mod q-1$. Since $c^{1-\sqrt{q}}=y_{0}^{\sqrt{q}}$,
$j(1-\sqrt{q})\equiv l(\sqrt{q}-1)\sqrt{q}\mod q-1.$
Hence
$j\equiv-l\sqrt{q}\mod\sqrt{q}+1.$ (21)
Since $\omega c^{\sqrt{q}}y_{0}=\omega_{0}^{k+j\sqrt{q}+l(\sqrt{q}-1)},$ $\operatorname{Tr}\omega c^{\sqrt{q}}y_{0}=0$ if and only if
$k+j\sqrt{q}+l(\sqrt{q}-1)\equiv\begin{cases}\frac{\sqrt{q}+1}{2}\mod\sqrt{q}+1&\text{ if $q$ is odd}\\ 0\mod\sqrt{q}+1&\text{ if $q$ is even}\end{cases}$ (22)
by Lemma 5.2. On the other hand,
$k+j\sqrt{q}+l(\sqrt{q}-1)\equiv k-l(q-\sqrt{q}+1)\mod\sqrt{q}+1$ (23)
by (21). Together with $\sqrt{q}\equiv-1\mod\sqrt{q}+1$ and $q\equiv-\sqrt{q}\equiv 1\mod\sqrt{q}+1,$ (22) and (23) imply
$k-3l\equiv\begin{cases}\frac{\sqrt{q}+1}{2}\mod\sqrt{q}+1&\text{ if $q$ is odd}\\ 0\mod\sqrt{q}+1&\text{ if $q$ is even}.\end{cases}$ (24)
Therefore the number of $y_{0}$’s for each of which (19) has $\mathbb{F}_{q}$-solutions, in fact $\sqrt{q}$-solutions, coincides with the number of $l$’s with $11\leq l\leq\sqrt{q}+$ 1 each of which satisfies (24) for the fixed $k$.
Let us consider the map
$\begin{array}[]{rccc}\varphi:&\mathbb{Z}/(\sqrt{q}+1)&\to&\mathbb{Z}/(\sqrt{q}+1)\\ &x&\mapsto&3x\end{array}.$
If $\sqrt{q}+1$ is not divisible by $3$, $\varphi$ is bijective. Hence there exists a unique $l$ satisfying (24) for any fixed $k$. If $3$ is a factor of $\sqrt{q}+1$, then
$|\varphi^{-1}(m)|=\begin{cases}3&\text{if $m\in\operatorname{Im}\varphi$}\\ 0&\text{else}.\end{cases}$
Hence if $k\equiv 0\mod 3$, there are three $l$’s satisfying (24), and otherwise no $l$’s. The condition $k\equiv 0\mod 3$ is equivalent to the condition $\omega^{\frac{\sqrt{q}+1}{3}}\in\mathbb{F}_{\sqrt{q}}^{\ast}.$ This completes the proof. ∎
Finally we pose a question: when $q\geq 16$, does a Hermitian-relative curve $C$ with $22\leq m_{C}\leq\sqrt{q}-$ 2 exist?
Acknowledgment. The second author was partially supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government(MSIT) (2022R1A2C1012291).