### v3

# Relatives of the Hermitian curve

Abstract
We introduce the notion of a relative of the Hermitian curve of degree $\sqrt{q}+1$ over $\mathbb{F}_{q}$, which is a plane curve defined by
$(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A{}^{t}\!(x,y,z)=0$
with $A\in GL(3,\mathbb{F}_{q})$, and study their basic properties, one of which is that the number of $\mathbb{F}_{q}$-points of any relative of the Hermitian curve of degree $\sqrt{q}+1$ is congruent to $1$ modulo $\sqrt{q}$.
In the latter part of this paper, we classify those curves having two or more rational inflexions.
Key Words: Plane curve, Finite field, Rational point
MSC: 14G15, 14H50, 14G05, 11G20
1 Introduction
Throughout this paper, $q$ denotes an even power of a prime number $p$. So $\sqrt{q}$ is also a power of $p$. For $A\in GL(3,\mathbb{F}_{q})$, $A^{\ast}$ denotes the matrix ${}^{t}\!A^{(\sqrt{q})}$, where the superscript $(\sqrt{q})$ indicates taking entry-wise $\sqrt{q}$-th power, and the superscripted prefix $t$ taking transpose.
The Hermitian curve is a plane curve defined by
$(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A\,{}^{t}\!(x,y,z)=0,$ (1)
with the condition $A=A^{\ast}$
In this case, we say that $A$ satisfies the Hermitian condition.
Forgetting the Hermitian condition, the equation (1) still defines a nonsingular plane curves of degree $\sqrt{q}+1$ over $\mathbb{F}_{q}$, which will be denoted by $C_{A}$. We call such a curve a relative of the Hermitian curve, or shortly, a Hermitian-relative curve, and want to study them.
All Hermitian-relative curves, including Hermitian curves themselves, are projectively equivalent each other over the algebraic closure $\overline{\mathbb{F}}_{q}$ of $\mathbb{F}_{q}$, which is due to Pardini [5, Proof of Proposition 3.7]. We will give a proof of this fact in the next section for the convenience of readers.
Any Hermitian-relative curve $C$ over $\mathbb{F}_{q}$ has a particular property on the number $N_{q}(C)$ of $\mathbb{F}_{q}$-points:
$N_{q}(C)\equiv 1\mod\sqrt{q}.$
The proof of this fact is also given in the following section. Since the equation (1) is homogeneous, the curve $C_{A}$ depends only on the image $\overline{A}$ of $A$ under the canonical map $GL(3,\mathbb{F}_{q})\to PGL(3,\mathbb{F}_{q})$. By abuse of notation, $\overline{A}$ will be frequently denoted by $A$, namely an element of $GL(3,\mathbb{F}_{q})$ and its image in $PGL(3,\mathbb{F}_{q})$ will be used interchangeably. So the equation $A=B$ in $PGL(3,\mathbb{F}_{q})$ means that there exists an element $\lambda\in\mathbb{F}_{q}^{\ast}$ so that $A=\lambda B$ in $GL(3,\mathbb{F}_{q})$.
Remark 1.1.
If $\overline{A}\in PGL(3,\mathbb{F}_{q})$ satisfies the Hermitian condition, that is to say, $\overline{A^{\ast}}=\overline{A}$, then we can choose an element $A\in GL(3,\mathbb{F}_{q})$ satisfying the Hermitian condition whose image in $PGL(3,\mathbb{F}_{q})$ is the assigned $\overline{A}$.
Proof.
Let $A$ be any source of $\overline{A}$. Then there is an element $\lambda\in\mathbb{F}_{q}^{\ast}$ such that $A^{\ast}=\lambda A$. Taking $\ast$-operation on the both side, $A=\lambda^{\sqrt{q}}A^{\ast}$. So $A=\lambda^{\sqrt{q}+1}A$, which implies $\lambda^{\sqrt{q}+1}=1$. Therefore there is an element $\rho\in\mathbb{F}_{q}$ so that $\rho^{1-\sqrt{q}}=\lambda$ (which is a special case of Hilbert’s Theorem 90). Then $\overline{\rho A}=\overline{A}$ and $\rho A$ satisfies the Hermitian condition. ∎
The polynomial of type (1) is a linear combination of the monomials
$x^{\sqrt{q}+1},y^{\sqrt{q}+1},z^{\sqrt{q}+1},x^{\sqrt{q}}y,y^{\sqrt{q}}z,z^{\sqrt{q}}x,xy^{\sqrt{q}},yz^{\sqrt{q}},zx^{\sqrt{q}}$
over $\mathbb{F}_{q}$, and vice versa. So those polynomials forms an $\mathbb{F}_{q}$-vector space of dimension $9$. Moreover the vector space is stable under the linear transformations on coordinates $x,y,z$. We should note that by a linear transformation
$\begin{pmatrix}x\\ y\\ z\end{pmatrix}\mapsto T^{-1}\begin{pmatrix}x\\ y\\ z\end{pmatrix}$
with $T\in PGL(3,\mathbb{F}_{q})$, the polynomial $(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A\begin{pmatrix}x\\ y\\ z\end{pmatrix}$ goes to
$(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})T^{\ast}AT\begin{pmatrix}x\\ y\\ z\end{pmatrix}.$
In other words, two relatives of the Hermitian curve $C_{A}$ and $C_{B}$, are projectively equivalent over $\mathbb{F}_{q}$ if and only if there is a certain matrix $T\in PGL(3,\mathbb{F}_{q})$ such that $B=T^{\ast}AT$ in $PGL(3,\mathbb{F}_{q})$. This is sometimes denoted by $A\sim B$.
Remark 1.2.
Even if $\det A=0$, the curve defined by (1) makes sense; however the curve is nonsingular if and only if $\det A\neq 0$. Actually, the partial derivatives of $f=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A\begin{pmatrix}x\\ y\\ z\end{pmatrix}$ are given by
$(f_{x},f_{y},f_{z})=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A.$ (2)
The properties mentioned above hold true after any field extension $\mathbb{F}_{q}$. As is explained in the next section, the tangent line at a point of $C_{A}$ contacts of multiplicity $\sqrt{q}$ in general, but at some special points of multiplicity $\sqrt{q}+1$. In the letter case, the point is called an inflexion. In Section 3, we give the classification up to $PGL(3,\mathbb{F}_{q})$ of the curves $C_{A}$ with two or more inflexions defined over $\mathbb{F}_{q}$. In the classification, three types appear and it can be summarized in Table 1.
Table 1: Classification
$N_{q}(C_{A})$ inflexions classes
(a) $\sqrt{q}^{3}+1$ $\sqrt{q}^{3}+1$ 1
(b) $\sqrt{q}+1$ $\sqrt{q}+1$ $\sqrt{q}$
(c) $q+1$ $2$ $\frac{1}{2}(\sqrt{q}+1)(\sqrt{q}-2)$
In the table, the first column indicates the label of each type which agrees with the item in Theorem 3.1, the second the number of $\mathbb{F}_{q}$-points of $C_{A}$ belongs the assigned type, the third the number of inflexions, and the fourth is the number of $\mathbb{F}_{q}$-equivalent classes in each type.
Annotation.
$\bullet$ Although a homogeneous coordinates of a point has an ambiguity by nonzero multiple, we understand all values of the coordinates of any $\mathbb{F}_{q}$-point are elements in $\mathbb{F}_{q}$. Moreover, we use two words “$\mathbb{F}_{q}$-point” and “rational point” interchangeably.
$\bullet$ The set of $\mathbb{F}_{q}$-points of a curve $C$ is denoted by $C(\mathbb{F}_{q})$.
$\bullet$ For a field, the multiplicative group of nonzero elements is marked by the superscript $\ast$.
$\bullet$ The trace map from $\mathbb{F}_{q}$ to $\mathbb{F}_{\sqrt{q}}$, that is, $\mathbb{F}_{q}\ni\alpha\mapsto\alpha+\alpha^{\sqrt{q}}\in\mathbb{F}_{\sqrt{q}},$ and the norm map $\mathbb{F}_{q}^{\ast}$ to $\mathbb{F}_{\sqrt{q}}^{\ast}$, that is, $\mathbb{F}_{q}^{\ast}\ni\alpha\mapsto\alpha^{\sqrt{q}+1}\in\mathbb{F}_{\sqrt{q}}^{\ast}$ are simply denoted by $\operatorname{Tr}$ and $\operatorname{Nm}$ respectively.
$\bullet$ Unless we specify the field of definition, ‘projectively equivalent’ means ‘projectively equivalent over $\mathbb{F}_{q}$’.
$\bullet$ Sometimes a curve defined by $g=0$ will be indicated by $\{g=0\}$.
2 Basic properties
We shall start this section with Pardini’s observation [5], which is our motivation to study the Hermitian-relative curve.
Theorem 2.1.
Let $f(x,y,z)\in\mathbb{F}_{q}[x,y,z]$ be a homogeneous polynomial of degree $\sqrt{q}+1$. The curve $C$ defined by $f(x,y,z)=0$ in $\mathbb{P}^{2}$ is projectively equivalent over $\overline{\mathbb{F}}_{q}$ to a Hermitian curve of degree $\sqrt{q}+1$ if and only if $C$ is a Hermitian-relative curve, that is, there exists $A\in PGL(3,\mathbb{F}_{q})$ such that
$f(x,y,z)=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A{}^{t}\!(x,y,z).$
Proof.
(Pardini) First suppose that $C$ is projectively equivalent to
$g(x,y,z)=x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+z^{\sqrt{q}+1}=0$
over $\overline{\mathbb{F}}_{q}$. Then there is a matrix $T=(t_{i,j})\in GL(3,\overline{\mathbb{F}}_{q})$ so that
$f(x,y,z)=g((x,y,z){}^{t}T),$
that is,
$f(x,y,z)$ $=\sum_{j=1}^{3}\left(t_{j1}x+t_{j2}y+t_{j3}z\right)^{\sqrt{q}+1}$
$=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})T^{\ast}T\begin{pmatrix}x\\ y\\ z\end{pmatrix}.$
Note that each entry of $T^{\ast}T$ gives a coefficient of distinct monomial of $f(x,y,z)$. So $T^{\ast}T\in PGL(3,\mathbb{F}_{q})$.
To see the converse, let us consider a morphism associated with an assigned $A\in PGL(3,\mathbb{F}_{q})$:
$\Phi_{A}:PGL(3,\overline{\mathbb{F}}_{q})\ni T\mapsto T^{\ast}AT\in PGL(3,\overline{\mathbb{F}}_{q}).$
If $\Phi_{A}(S)=\Phi_{A}(T)$ for $S,T\in PGL(3,\overline{\mathbb{F}}_{q})$, then
$(ST^{-1})^{\ast}A(ST^{-1})=A$
which means $ST^{-1}$ gives a linear automorphism of $C_{A}$ over $\overline{\mathbb{F}}_{q}$ in $\mathbb{P}^{2}$.
The automorphism group of $C_{A}$ over $\overline{\mathbb{F}}_{q}$ is finite if $\sqrt{q}>2$, because the genus of $C_{A}$ is grater than $1$. When $\sqrt{q}=2$, $C_{A}$ is an elliptic curve over $\overline{\mathbb{F}}_{q}$ in $\mathbb{P}^{2}$. The linear automorphism group of any elliptic curve in $\mathbb{P}^{2}$ is also finite. Therefore the image of $\Phi_{A}$ is a constructible set (which is a theorem of Chevalley [4, Corollary 2 in I §8]) of dimension $8$ in the target space $PGL(3,\overline{\mathbb{F}}_{q})$. Therefore two images of $\Phi_{A}$ and $\Phi_{B}$ meet in $PGL(3,\overline{\mathbb{F}}_{q})$ for any two $A,B\in PGL(3,\mathbb{F}_{q})$, that is, there is an element $T\in PGL(3,\overline{\mathbb{F}}_{q})$ such that $T^{\ast}AT=B$. In particular, if one takes as $B=I_{3}$, one gets a projective transformation over $\overline{\mathbb{F}}_{q}$ which sends $C_{A}$ to a Hermitian curve. ∎
In the next corollary, $T_{P}(C)$ denotes the (embedded) tangent line to a plane curve $C$ at $P$, and $i(C.T_{P}(C);P)$ the intersection multiplicity of $C$ and $T_{P}(C)$ at $P$.
Corollary 2.2.
Suppose that $q>4$. Among nonsingular plane curves of degree $\sqrt{q}+1$, a Hermitian-relative curve $C$ is characterized by the property: for any $\overline{\mathbb{F}}_{q}$-point $P$ of $C$, $i(C.T_{P}(C);P)$ is either $\sqrt{q}$ or $\sqrt{q}+1$.
Proof.
This characterization comes from [2, Theorem 6.1, and Theorem 3.5] together with Theorem 2.1
The following is the main purpose of this section.
Theorem 2.3.
Let $C$ be a relative of the Hermitian curve of degree $\sqrt{q}+1$ over $\mathbb{F}_{q}$. Then
$N_{q}(C)\equiv 1\mod\sqrt{q}.$
In particular, the set $C(\mathbb{F}_{q})$ of $\mathbb{F}_{q}$-points is not empty.
Proof.
Let $g=\frac{1}{2}\sqrt{q}(\sqrt{q}-1)$, which is the genus of any Hermitian-relative curve. Let $C^{\prime}$ be a Hermitian curve of degree $\sqrt{q}+1$, e.g., $C^{\prime}$ is one defined by $x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+z^{\sqrt{q}+1}=0.$ By Theorem 2.1, $C$ and $C^{\prime}$ are projectively equivalent over some extension field $\mathbb{F}_{q^{s}}$. Let
$Z_{C}(t)=\prod_{j=1}^{2g}(1-\alpha_{j}t)/(1-t)(1-qt)$
be the zeta function of $C$ over $\mathbb{F}_{q}$. By Weil’s theorem,
$N_{q^{s}}(C)=1+q^{s}-\sum_{j=1}^{2g}\alpha_{j}^{s}\text{\ and \ }\alpha_{j}=\sqrt{q}\eta_{j}\text{\ with \ }|\eta_{j}|=1.$ (3)
On the other hand, since $C^{\prime}$ is a maximal curve over $\mathbb{F}_{q}$, its zeta function is
$Z_{C^{\prime}}(t)=\left(1-(-\sqrt{q})t\right)^{2g}/(1-t)(1-qt).$
Since $C$ and $C^{\prime}$ are isomorphic over $\mathbb{F}_{q^{s}}$,
$\{\alpha_{1}^{s},\dots,\alpha_{2g}^{s}\}=\{\overbrace{(-\sqrt{q})^{s},\dots,(-\sqrt{q})^{s}}^{2g}\}.$
(As for this classical theory, one may consult [7, Chapter 5].)
Since we may choose $s$ to be even, $\eta_{j}^{s}=1$ for each $j=1,\dots,2g$. In particular, $\eta_{j}$ is an algebraic integer, and so is $\sum_{j=1}^{2g}\eta_{j}$. Going back to the formula (3) (with $s=1$),
$\sum_{j=1}^{2g}\eta_{j}=\frac{q+1-N_{q}(C)}{\sqrt{q}}.$
Since $\sqrt{q}$ is a rational integer by our assumption, $\sum_{j=1}^{2g}\eta_{j}$ is a rational number. Therefore it is a rational integer. Hence
$N_{q}(C)=q+1-\sqrt{q}(\sum_{j=1}^{2g}\eta_{j})\equiv 1\mod\sqrt{q}.$
This completes the proof. ∎
Corollary 2.4.
For a relative of the Hermitian curve $C$ over $\mathbb{F}_{q}$, there is a nonnegative integer $m=m_{C}$ with
$m=q\text{\ or \ }\sqrt{q}+2\geq m\geq 0$
so that $N_{q}(C)=m\sqrt{q}+1$. Moreover $m_{C}=q$ if and only if either $C$ is Hermitian, or $q=4$ and $C$ is projectively equivalent to the curve
$x^{\sqrt{q}+1}+\omega y^{\sqrt{q}+1}+\omega^{2}z^{\sqrt{q}+1}=0.$
over $\mathbb{F}_{4}$, where $\omega\in\mathbb{F}_{4}\setminus\mathbb{F}_{2}$.
Proof.
The additional statement is somewhat classical [1, 6, 3]. For the first statement, when $q=4$, there is nothing new from Theorem 2.3. Hence we assume that $q>4$. Let $C=C_{A}$, and $(x_{0},y_{0},z_{0})\in C(\mathbb{F}_{q})$. Then $(x_{0}^{\sqrt{q}},y_{0}^{\sqrt{q}},z_{0}^{\sqrt{q}})A{}^{t}\!(x_{0},y_{0},z_{0})=0.$ Taking $\sqrt{q}$-th power, and noting $x_{0}^{q}=x_{0},y_{0}^{q}=y_{0},z_{0}^{q}=z_{0}$, we know that $(x_{0},y_{0},z_{0})\in C_{A^{\ast}}$. If $C$ is not Hermitian, $C_{A}$ and $C_{A^{\ast}}$ are distinct irreducible curves. Hence the intersection of these two curves consists of at most $(\sqrt{q}+1)^{2}$ points, which gives an upper bound for $N_{q}(C)$. ∎
Before closing this section, we shall digress a little to explain a couple of geometric properties of relatives of the Hermitian curve.
Proposition 2.5.
For $A\in GL(3,\mathbb{F}_{q})$, the dual curve of $C_{A}$ is $C_{{}^{t}\!A^{-1}}$.
Proof.
Let $u,v,w$ be the dual coordinate of the dual projective plane with point-line incidence $ux+vy+wz=0$. By (2), the equation after eliminating $x,y,z$ from
$(u,v,w)=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A$ (4)
and (1) is an equation of the dual curve. Taking entry-wise $\sqrt{q}$-th power of (1),
$(x^{q},y^{q},z^{q})A^{(\sqrt{q})}\begin{pmatrix}x^{\sqrt{q}}\\ y^{\sqrt{q}}\\ z^{\sqrt{q}}\end{pmatrix}=0,$
which can be written
$(u^{\sqrt{q}},v^{\sqrt{q}},w^{\sqrt{q}}){}^{t}\!A^{-1}\begin{pmatrix}u\\ v\\ w\end{pmatrix}=0$
by (4). This completes the proof. ∎
For a Hermitian-relative curve $C_{A}$, the curve $C_{A^{\ast}}$ is called the mirror curve of $C_{A}$. Of course, the mirror curve of $C_{A^{\ast}}$ is the original one $C_{A}$, and $C_{A}$ is Hermitian if and only if $C_{A^{\ast}}=C_{A}$.
Proposition 2.6.
For a mirror pair $(C_{A},C_{A^{\ast}})$, the following properties hold.
  1. (1)
    $C_{A}(\mathbb{F}_{q})=C_{A^{\ast}}(\mathbb{F}_{q})$.
  2. (2)
    For $P\in C_{A}(\mathbb{F}_{q})=C_{A^{\ast}}(\mathbb{F}_{q})$, $P$ is an inflexion of $C_{A}$ if and only if an inflexion of $C_{A^{\ast}}$. In this case, $T_{P}(C_{A})=T_{P}(C_{A}^{\ast})$.
  3. (3)
    For a non-inflexion $P\in C_{A}(\mathbb{F}_{q})$, then there is a point $P^{\prime}\in C_{A}(\mathbb{F}_{q})$ so that $C_{A}.T_{P}C_{A}=\sqrt{q}P+P^{\prime}$. In this case, $C_{A^{\ast}}.T_{P^{\prime}}C_{A^{\ast}}=\sqrt{q}P^{\prime}+P$. In particular $P^{\prime}$ is a non-inflexion of $C_{A^{\ast}}$.
Proof.
If $C_{A}$ is Hermitian, then the assertions (1) and (2) are trivially true, and (3) is nonsense. So $C_{A}$ may be assumed to be non-Hermitian.
We already saw the assertion (1) in the proof of Corollary 2.4.
(2) Since $(T^{\ast}AT)^{\ast}=T^{\ast}A^{\ast}T$ for any $T\in PGL(3,\mathbb{F}_{q})$, we may assume that $P=(0,0,1)$ and $T_{P}C_{A}=\{y=0\}$. Let $A=(a_{ij})$. Since $P\in C_{A}$, $a_{33}=0$. Hence an affine equation of $C_{A}$ on $\{z\neq 0\}$ is given by
$a_{11}x^{\sqrt{q}+1}+(a_{12}y+a_{13})x^{\sqrt{q}}+(a_{21}x+a_{22}y+a_{23})y^{\sqrt{q}}+a_{31}x+a_{32}y=0.$
Since $T_{P}C_{A}=\{y=0\}$ and the origin $P$ is an inflexion of $C_{A}$, we have $a_{31}=0$, $a_{13}=0$ and $a_{11}\neq 0$. Hence we may assume that $a_{11}=1$. To sum up,
$A=\begin{pmatrix}1&a_{12}&0\\ a_{21}&a_{22}&a_{23}\\ 0&a_{32}&0\end{pmatrix}.$
Hence
$A^{\ast}=\begin{pmatrix}1&a_{21}^{\sqrt{q}}&0\\ a_{12}^{\sqrt{q}}&a_{22}^{\sqrt{q}}&a_{32}^{\sqrt{q}}\\ 0&a_{23}^{\sqrt{q}}&0\end{pmatrix},$
in other words, an affine equation of $C_{A^{\ast}}$ is
$x^{\sqrt{q}+1}+a_{21}^{\sqrt{q}}yx^{\sqrt{q}}+(a_{12}^{\sqrt{q}}x+a_{22}^{\sqrt{q}}y+a_{32}^{\sqrt{q}})y^{\sqrt{q}}+a_{32}^{\sqrt{q}}y=0.$
This shows that the origin is also an inflexion of $C_{A^{\ast}}$ with tangent line $\{y=0\}$.
(3) By Corollary 2.2, $C_{A}.T_{P}C_{A}=\sqrt{q}P+P^{\prime}$ for some point $P^{\prime}$. Since two divisors $C_{A}.T_{P}C_{A}$ and $\sqrt{q}P$ are defined over $\mathbb{F}_{q}$, so is the residual one $P^{\prime}$. Hence $P^{\prime}\in C_{A}(\mathbb{F}_{q})=C_{A^{\ast}}(\mathbb{F}_{q})$. Let $P=(x_{0},y_{0},z_{0})$ and $P^{\prime}=(x_{1},y_{1},z_{1})$. Then
$(x_{0}^{\sqrt{q}},y_{0}^{\sqrt{q}},z_{0}^{\sqrt{q}})A{}^{t}\!(x_{1},y_{1},z_{1})$
by Remark 2. Hence
$(x_{1}^{\sqrt{q}},y_{1}^{\sqrt{q}},z_{1}^{\sqrt{q}})A^{\ast}{}^{t}\!(x_{0},y_{0},z_{0}),$
which means $P\in C_{A^{\ast}}.T_{P^{\prime}}C_{A^{\ast}}$, and hence $C_{A^{\ast}}.T_{P^{\prime}}C_{A^{\ast}}=\sqrt{q}P^{\prime}+P.$
3 Hermitian-relative curves with two or more rational inflexions
The purpose of this section is to give the classification of the Hermitian-relative curves having two or more rational inflexions, which is summarized in the following theorem.
Theorem 3.1.
Let $C$ be a relative of the Hermitian curve. Suppose that $C$ has at least two rational inflexions. Then $C$ is projectively equivalent to
$x^{\sqrt{q}}y+\omega xy^{\sqrt{q}}+z^{\sqrt{q}+1}=0$ (5)
[…]
  1. (a)
[…]
  2. (b)
[…]
    $x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+\eta z^{\sqrt{q}+1}=0$ (6)
[…]
    For two curves $x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+\eta z^{\sqrt{q}+1}=0$ and $x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+\eta^{\prime}z^{\sqrt{q}+1}=0$ with $\eta,\,\eta^{\prime}\in\mathbb{F}_{q}^{\ast}\setminus\mathbb{F}_{\sqrt{q}}^{\ast}.$ are projectively equivalent to each other if and only if $\eta\eta^{\prime-1}\in\mathbb{F}_{\sqrt{q}}^{\ast}$. Therefore, there are exactly $\sqrt{q}$ non-equivalent Hermitian-relative curves, each of which has $\sqrt{q}+1$ rational points and all of them are inflexions.
  3. (c)
[…]
3.1 Some equations
Since any trace map and norm map between finite fields are surjective, it is easy to show Hilbert’s Theorem 90 in our case:
Lemma 3.2.
There are two exact sequences:
$\begin{array}[]{rcccccccl}0\to&\mathbb{F}_{\sqrt{q}}&\stackrel{{\scriptstyle i}}{{\to}}&\mathbb{F}_{q}&\to&\mathbb{F}_{q}&\xrightarrow{\operatorname{Tr}}&\mathbb{F}_{\sqrt{q}}&\to 0\\ &&&\alpha&\mapsto&\alpha-\alpha^{\sqrt{q}}&&&\end{array}$ (7)
$\begin{array}[]{rcccccccl}1\to&\mathbb{F}_{\sqrt{q}}^{\ast}&\stackrel{{\scriptstyle i}}{{\to}}&\mathbb{F}_{q}^{\ast}&\to&\mathbb{F}_{q}^{\ast}&\xrightarrow{\operatorname{Nm}}&\mathbb{F}_{\sqrt{q}}^{\ast}&\to 1\\ &&&\alpha&\mapsto&\alpha^{1-\sqrt{q}}&&&\end{array},$ (8)
where $i$ means natural inclusion.
Corollary 3.3.
  1. (i)
    The equation of $X$ with $\beta\in\mathbb{F}_{q}$
    $X^{\sqrt{q}}-X-\beta=0$ (9)
[…]
  2. (ii)
    The equation of $X$ with $\beta\in\mathbb{F}_{q}^{\ast}$
    $X^{\sqrt{q}-1}-\beta=0$ (10)
[…]
Proof.
(i) Let $\alpha\in\mathbb{F}_{q}$ be a root of (9). Then $\beta=\alpha^{\sqrt{q}}-\alpha$. Hence $\operatorname{Tr}\beta=0$. Conversely, if $\operatorname{Tr}\beta=0$, there is an element $\alpha\in\mathbb{F}_{q}$ such that $\alpha^{\sqrt{q}}-\alpha=\beta$ by (7). When $\alpha$ is a root of (9), the set of all roots is $\{\alpha+\gamma|\gamma\in\mathbb{F}_{\sqrt{q}}\}$.
(ii) can be proved by using (8). ∎
The above corollary can be generalized slightly as follows:
Proposition 3.4.
Let $\alpha\in\mathbb{F}_{q}^{\ast}$ and $\beta\in\mathbb{F}_{q}$, and consider the equation in $X$:
$X^{\sqrt{q}}+\alpha X+\beta=0.$ (11)
  1. (i)
[…]
  2. (ii)
[…]
Proof.
For a fixed $\alpha\in\mathbb{F}_{q}^{\ast}$, consider the $\mathbb{F}_{\sqrt{q}}$-linear map $\varphi_{\alpha}:\mathbb{F}_{q}\to\mathbb{F}_{q}$ given by $\varphi_{\alpha}(z)=z^{\sqrt{q}}+\alpha z$. Since $\operatorname{Nm}(-\alpha)=\operatorname{Nm}\alpha$, the kernel of $\varphi_{\alpha}$ is nontrivial if and only if $\operatorname{Nm}\alpha=1$ by Corollary 3.3 (ii).
Therefore if $\operatorname{Nm}\alpha\neq 1$, then $\varphi_{\alpha}$ is an isomorphism. Hence there is a unique element of $\mathbb{F}_{q}$, say $z_{0}$, so that $\varphi_{\alpha}(z_{0})=-\beta$. So the proof of (ii) is done.
Next suppose that $\operatorname{Nm}\alpha=1$. In this case, $\dim\operatorname{Ker}\varphi_{\alpha}>0$ from the first paragraph of this proof, in particular there is an element $z_{0}\in\mathbb{F}_{q}^{\ast}$ such that $z_{0}^{\sqrt{q}-1}+\alpha=0$. Hence
$\operatorname{Ker}\varphi_{\alpha}=\{\zeta z_{0}\mid\zeta\in\mathbb{F}_{\sqrt{q}}^{\ast}\}\cup\{0\}.$
If (11) has a solution $z_{1}$, the set of all solutions of (11) is
$\{z_{1}+\zeta z_{0}\mid\zeta\in\mathbb{F}_{\sqrt{q}}^{\ast}\}\cup\{z_{1}\},$
which is a subset of $\mathbb{F}_{q}$. So the additional statement of (i) is done.
Now we prove the first part of (i). To prove the ‘only if’ part, we may suppose that $\beta\neq 0$. Let $z_{1}$ be an $\mathbb{F}_{q}$-solution of (11). Then $\alpha=-\frac{z_{1}^{\sqrt{q}}+\beta}{z_{1}}$. Since $\operatorname{Nm}\alpha=1$ and $(z_{1}^{\sqrt{q}})^{\sqrt{q}}=z_{1}$, we have
11 11 $=$ $\operatorname{Nm}\alpha=-(\frac{z_{1}^{\sqrt{q}}+\beta}{z_{1}})^{\sqrt{q}}(-\frac{z_{1}^{\sqrt{q}}+\beta}{z_{1}})$
$=$ $\frac{z_{1}^{\sqrt{q}+1}+\beta^{\sqrt{q}}z_{1}^{\sqrt{q}}+\beta z_{1}+\beta^{\sqrt{q}+1}}{z_{1}^{\sqrt{q}+1}}.$
Hence
$z_{1}^{\sqrt{q}}+\beta^{1-\sqrt{q}}z_{1}+\beta=0.$
Since $z_{1}$ satisfies (11), we get $\alpha=\beta^{1-\sqrt{q}}$. Conversely, if $\beta=0$, obviously (11) has an $\mathbb{F}_{q}$-solution. Hence we may consider the case $\alpha=\beta^{1-\sqrt{q}}$ with $\beta\neq 0$. In this case, (11) is
$X^{\sqrt{q}}+\beta^{1-\sqrt{q}}X+\beta=0$ (12)
which is equivalent to
$(\beta X)^{\sqrt{q}}+(\beta X)+\beta^{\sqrt{q}+1}=0.$ (13)
Since
$(-\beta^{\sqrt{q}+1})^{\sqrt{q}-1}=\beta^{q-1}=1$
because $\beta\in\mathbb{F}_{q}^{\ast}$, we know $-\beta^{\sqrt{q}+1}\in\mathbb{F}_{\sqrt{q}}^{\ast}$. Since $\operatorname{Tr}:\mathbb{F}_{q}\to\mathbb{F}_{\sqrt{q}}$ is surjective, the equation (13) in $(\beta X)$ has a solution in $\mathbb{F}_{q}$, and also (12) has. ∎
3.2 Proof of Theorem 3.1
Lemma 3.5.
Let $C$ be a relative of the Hermitian curve with at least two rational inflexions. Then we can choose a system of coordinates $x,y,z$ of $\mathbb{P}^{2}$ over $\mathbb{F}_{q}$ such that
  • $P=(1,0,0)$ and $Q=(0,1,0)$ are inflexions of $C$, and
  • $T_{P}(C)=\{y=0\}$ and $T_{Q}(C)=\{x=0\}$.
Proof.
Since $C.T_{P}(C)=(\sqrt{q}+1)P$ and $C.T_{Q}(C)=(\sqrt{q}+1)Q$, three $\mathbb{F}_{q}$-lines $T_{P}(C)$, $T_{Q}(C)$ and $PQ$ are not concurrent, where $PQ$ denotes the line through $P$ and $Q$. Therefore, we may choose coordinates as $T_{P}(C)=\{y=0\}$, $T_{Q}(C)=\{x=0\}$ and $PQ=\{z=0\}$. ∎
Lemma 3.6.
Under the situation described in Lemma 3.5, the curve is defined by $(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A\,{}^{t}\!(x,y,z)=0,$ where
$A=\begin{pmatrix}0&a_{12}&0\\ a_{21}&0&0\\ 0&0&a_{33}\end{pmatrix}\in PGL(3,\mathbb{F}_{q})\ \text{with}\ a_{12}a_{21}a_{33}\neq 0.$
In other words, $C$ is defined by
$a_{12}x^{\sqrt{q}}y+a_{21}xy^{\sqrt{q}}+a_{33}z^{\sqrt{q}+1}=0.$ (14)
Proof.
Let $C=C_{A}$, where $A=(a_{ij})\in PGL(3,\mathbb{F}_{q})$. Since $(1,0,0)A{}^{t}\!(1,0,0)=0$ and $(0,1,0)A{}^{t}\!(0,1,0)=0$, $a_{11}=a_{22}=0$.
Note that the tangent line at $(b_{1},b_{2},b_{3})\in C$ is given by
$(b_{1}^{\sqrt{q}},b_{2}^{\sqrt{q}},b_{3}^{\sqrt{q}})A{}^{t}\!(x,y,z)=0.$
Hence $a_{13}=0$ because $T_{P}(C)=\{y=0\}$, and $a_{23}=0$ because $T_{Q}(C)=\{x=0\}$. Therefore an equation of $C$ is of the form
$a_{12}x^{\sqrt{q}}y+a_{21}xy^{\sqrt{q}}+a_{31}xz^{\sqrt{q}}+a_{32}yz^{\sqrt{q}}+a_{33}z^{\sqrt{q}+1}=0.$
Since $P$ is an inflexion of $C$ and $T_{P}(C)=\{y=0\}$, we have $C.\{y=0\}=(\sqrt{q}+1)P$. Hence the equation
$a_{31}xz^{\sqrt{q}}+a_{33}z^{\sqrt{q}+1}=0$
has the root at $z=0$ of multiplicity $(\sqrt{q}+1)$. So $a_{31}=0$. Similarly, since $C.\{x=0\}=(\sqrt{q}+1)Q$, from the equation $a_{32}yz^{\sqrt{q}}+a_{33}z^{\sqrt{q}+1}=0,$ we know $a_{32}=0$. Finally $\det A\neq 0$ implies $a_{12}a_{21}a_{33}\neq 0$. ∎
Corollary 3.7.
After a suitable projective transformation over $\mathbb{F}_{q}$, $C$ is defined by
$x^{\sqrt{q}}y+\omega xy^{\sqrt{q}}+z^{\sqrt{q}+1}=0$
with $\omega\in\mathbb{F}_{q}^{\ast}$. Conversely, a curve defined by this type of equation is Hermitian-relative with two or more inflexions.
Proof.
Since $a_{33}\neq 0$, we may suppose that $a_{33}=1$ in (14). If one consider $a_{12}y$ as new $y$, then the equation becomes
$x^{\sqrt{q}}y+a_{21}a_{12}^{-{\sqrt{q}}}xy^{\sqrt{q}}+z^{\sqrt{q}+1}=0.$
Put $a_{21}a_{12}^{-{\sqrt{q}}}=\omega.$
To confirm the last assertion is not difficult, actually $(1,0,0)$ and $(0,1,0)$ are inflexions. ∎
By Corollary 3.7, we have done the proof of the first paragraph of Theorem 3.1.
Now we investigate the curve (5), hereafter, this curve will be denoted by $C_{\omega}$.
Lemma 3.8.
The number of $\mathbb{F}_{q}$-points on $C_{\omega}\cap\{z=0\}$ is $\sqrt{q}+1$ if $\operatorname{Nm}\omega=1$, and $2$ if $\operatorname{Nm}\omega\neq 1$.
Proof.
For any $\omega$, $P=(1,0,0)$ and $Q=(0,1,0)$ are rational inflexions on $C_{\omega}$. Since $\{z=0\}\cap\{y=0\}=\{P\}$, it is enough to count the number of $\mathbb{F}_{q}$-solutions of (5) when $y=1$ and $z=0$. The equation in question is $x^{\sqrt{q}}+\omega x=0.$ Except for $Q$, the other points of $C_{\omega}\cap\{z=0\}\cap\{y=1\}$ correspond to the roots of $x^{\sqrt{q}-1}+\omega=0.$ By Corollary 3.3, the number of $\mathbb{F}_{q}$-solutions of this equation is
$\left\{\begin{array}[]{cl}\sqrt{q}-1&\text{if $\operatorname{Nm}\omega\,(=\operatorname{Nm}(-\omega))=1$}\\ 0&\text{ else.}\end{array}\right.$
Adding $2$ coming from $\{P,Q\}$ in each cases, we have the desired numbers. ∎
Next we compute the number of rational points of $C_{\omega}$ outside of $\{z=0\}$.
Lemma 3.9.
The number of $\mathbb{F}_{q}$-solutions of affine equation
$x^{\sqrt{q}}y+\omega xy^{\sqrt{q}}+1=0$ (15)
is
$\left\{\begin{array}[]{cl}\sqrt{q}(q-1)&\text{ if $\omega=1$}\\ 0&\text{ if $\operatorname{Nm}\omega=1$ and $\omega\neq 1$}\\ q-1&\text{ if $\operatorname{Nm}\omega\neq 1$}.\end{array}\right.$
Proof.
If the affine equation (15) has an $\mathbb{F}_{q}$-solution $(x_{0},y_{0})$, then $x_{0}y_{0}^{\sqrt{q}}$ is a root of the equation in $T$
$T^{\sqrt{q}}+\omega T+1=0.$ (16)
Conversely if $t_{0}$ is a root of (16), then $(\frac{t_{0}}{y_{0}^{\sqrt{q}}},y_{0})$ is a solution of (15) for any $y_{0}\in\mathbb{F}_{q}^{\ast}$. Therefore the number of solutions of (15) is
[…]
Suppose that $\operatorname{Nm}\omega=1$. Then, by Proposition 3.4 (i), (16) has an $\mathbb{F}_{q}$-solution if and only if $\omega=1$. In this case, it has exactly $\sqrt{q}$ rational solutions, and totally (15) has $\sqrt{q}(q-1)$ rational solutions.
Suppose that $\operatorname{Nm}\omega\neq 1$. Then, from the second part of the same proposition, (16) has a unique $\mathbb{F}_{q}$-solution, and then (15) has $(q-1)$ rational solutions. ∎
The following three lemmas will be used for showing the part (b) of Theorem 3.1.
The element of $PGL(3,\mathbb{F}_{q})$ which comes from the diagonal matrix with entries $a,b,c$ from the upper left to the lower right will be denoted by $\operatorname{diag}[a,b.c]$.
Lemma 3.10.
For $\eta\in\mathbb{F}_{q}^{\ast}$, we have
$\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&\eta\end{pmatrix}\sim\operatorname{diag}[1,1,\eta].$
Proof.
Choose $u\in\mathbb{F}_{q}$ so that $\operatorname{Tr}u=1$ and $\alpha\in\mathbb{F}_{q}^{\ast}$ so that $\operatorname{Nm}\alpha=-1$. Let $T_{0}=\begin{pmatrix}u&\alpha u^{\sqrt{q}}&0\\ 1&-\alpha&0\\ 0&0&1\end{pmatrix}.$ Then by straightforward computation, we know that $T_{0}^{\ast}\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&\omega\end{pmatrix}T_{0}=\operatorname{diag}[1,1,\eta].$
Lemma 3.11.
Let $\eta\in\mathbb{F}_{q}^{\ast}\setminus\mathbb{F}_{\sqrt{q}}^{\ast}$. Then the number of rational points of $C_{\operatorname{diag}[1,1,\eta]}$ is $\sqrt{q}+1$, and these rational points are on the line $\{z=0\}$.
Proof.
The curve $C_{\operatorname{diag}[1,1,\eta]}$ is defined by
$x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+\eta z^{\sqrt{q}+1}=0.$ (17)
If $(x_{0},y_{0},z_{0})\in(\mathbb{F}_{q})^{3}$ is a solution of (17), then it is a solution of the simultaneous equations:
[…]
because $x_{0}^{\sqrt{q}+1},y_{0}^{\sqrt{q}+1},z_{0}^{\sqrt{q}+1}$ are elements of $\mathbb{F}_{\sqrt{q}}$, and $\{1,\eta\}$ are linearly independent over $\mathbb{F}_{\sqrt{q}}$. Therefore
$C_{\operatorname{diag}[1,1,\eta]}(\mathbb{F}_{q})=\{(\zeta,1,0)|\operatorname{Nm}\zeta=-1\}\subset\{z=0\}.$
This completes the proof. ∎
Lemma 3.12.
Let $\eta,\eta^{\prime}\in\mathbb{F}_{q}^{\ast}\setminus\mathbb{F}_{\sqrt{q}}^{\ast}$. Then there is an element $T\in PGL(3,\mathbb{F}_{q})$ such that
$T^{\ast}\operatorname{diag}[1,1,\eta]T=\operatorname{diag}[1,1,\eta^{\prime}]$
if and only if $\eta^{\prime}\eta^{-1}\in\mathbb{F}_{\sqrt{q}}^{\ast}$.
Proof.
Let $\lambda=\eta^{\prime}\eta^{-1}$, and suppose that $\lambda\in\mathbb{F}_{\sqrt{q}}^{\ast}$. Since $\operatorname{Nm}$ is surjective, there is an element $\alpha\in\mathbb{F}_{q}^{\ast}$ such that $\alpha^{\sqrt{q}+1}=\lambda$. So
$\operatorname{diag}[1,1,\alpha]^{\ast}\operatorname{diag}[1,1,\eta]\operatorname{diag}[1,1,\alpha]=\operatorname{diag}[1,1,\eta^{\prime}].$
Conversely, let $T=(a_{ij})\in GL(3,\mathbb{F}_{q})$ such that
$T^{\ast}\operatorname{diag}[1,1,\eta]T=\rho\operatorname{diag}[1,1,\eta^{\prime}],$
where we are regarding matrices as in $GL(3,\mathbb{F}_{q})$ and $\rho\in\mathbb{F}_{q}^{\ast}$. For both curves $C_{\operatorname{diag}[1,1,\eta]}$ and $C_{\operatorname{diag}[1,1,\eta^{\prime}]}$, their sets of $\mathbb{F}_{q}$-points consist of $\sqrt{q}+1$ points on the line $\{z=0\}$ by Lemma 3.11. Hence the line $\{z=0\}$ is stable by the linear transformation of $\mathbb{P}^{2}$ corresponding $T$, that is
$\begin{pmatrix}a_{11}&a_{12}&a_{13}\\ a_{21}&a_{22}&a_{23}\\ a_{31}&a_{32}&a_{33}\end{pmatrix}\begin{pmatrix}\ast\\ \ast\\ 0\end{pmatrix}=\begin{pmatrix}\ast\\ \ast\\ 0\end{pmatrix}.$
Hence $a_{31}=a_{32}=0$, and
$\begin{pmatrix}a_{11}^{\sqrt{q}}&a_{21}^{\sqrt{q}}&0\\ a_{12}^{\sqrt{q}}&a_{22}^{\sqrt{q}}&0\\ a_{13}^{\sqrt{q}}&a_{23}^{\sqrt{q}}&a_{33}^{\sqrt{q}}\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&\eta\end{pmatrix}\begin{pmatrix}a_{11}&a_{12}&a_{13}\\ a_{21}&a_{22}&a_{23}\\ 0&0&a_{33}\end{pmatrix}=\rho\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&\eta^{\prime}\end{pmatrix}.$
Comparing $(1,1)$-component of the both sides, we have $a_{11}^{\sqrt{q}+1}+a_{21}^{\sqrt{q}+1}=\rho$, which guarantees that $\rho$ is an element of $\mathbb{F}_{\sqrt{q}}^{\ast}$. Taking determinant of the both sides, we have
$\det T^{(\sqrt{q})}\cdot\eta\cdot\det T=\rho^{3}\eta^{\prime}.$
Hence $\eta^{\prime}\eta^{-1}=(\det T)^{\sqrt{q}+1}\rho^{-3}\in\mathbb{F}_{\sqrt{q}}^{\ast}$
Proof of Theorem 3.1.
(a) Suppose that $\omega=1$. Then by Lemmas 3.8, 3.9, the curve $C_{\omega}$ has $\sqrt{q}^{3}+1$ rational points. Hence it is Hermitian if $q\neq 4$ [1], and all rational points on a Hermitian curve are inflexions. When $q=4$, a non-Hermitian curve with $9$ rational points can’t have inflexions [3, Remark 3.10].
(b) Suppose that $\operatorname{Nm}\omega=1$ and $\omega\neq 1$. Then by Lemmas 3.8, 3.9, the curve $C_{\omega}$ has $\sqrt{q}+1$ rational points, and those points lie on the line $\{z=0\}$. Let $R$ be one of those points. If $R$ is not an inflexion, then $C_{\omega}.T_{R}(C_{\omega})=\sqrt{q}R+R^{\prime}$ with $R^{\prime}\neq R$. Since $R^{\prime}$ is $\mathbb{F}_{q}$-point too, $R^{\prime}$ lies on the line $\{z=0\}$. Hence $T_{R}(C_{\omega})$ coincides with the line $\{z=0\}$, which is impossible. Therefore all the $\sqrt{q}+1$ rational points are inflexions.
Next we show that $C_{\omega}$ with $\operatorname{Nm}\omega=1$ and $\omega\neq 1$ is projectively equivalent to
$x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+\eta z^{\sqrt{q}+1}=0$ (18)
for a certain $\eta\in\mathbb{F}_{q}^{\ast}\setminus\mathbb{F}_{\sqrt{q}}^{\ast}.$ Since $\operatorname{Nm}\omega=1$, there is an element $\beta\in\mathbb{F}_{q}^{\ast}\setminus\mathbb{F}_{\sqrt{q}}^{\ast}$ such that $\beta^{1-\sqrt{q}}=\omega$. Using this $\beta$, (5) can be rewritten as
$(\beta x)^{\sqrt{q}}y+(\beta x)y^{\sqrt{q}}+\beta^{\sqrt{q}}z^{\sqrt{q}+1}=0.$ (19)
Replacing $\beta x$ with the new $x$ and applying Lemma 3.10, we get the equation
$x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+\beta^{\sqrt{q}}z^{\sqrt{q}+1}=0.$
Replace $\beta^{\sqrt{q}}$ by $\eta$, then we have (18). Conversely, tracing the reverse process, we know that the curve defined by (18) is projectively equivalent to $C_{\eta^{\sqrt{q}-1}}$.
The last statement in (b) is just Lemma 3.12.
(c) Suppose that $\operatorname{Nm}\omega\neq 1$. From Lemmas 3.8, 3.9, $C_{\omega}(\mathbb{F}_{q})$ consists of $q+1$ rational points. Note that $P=(1,0,0)$ and $Q=(0,1,0)$ are inflexions, and they are only $\mathbb{F}_{q}$-points lie on the line $\{z=0\}$ by Lemma 3.5.
We will show that any point $R\in C_{\omega}(\mathbb{F}_{q})\setminus\{P,Q\}$ is not an inflexion. Since $R$ does not lie on the line $\{z=0\}$, we may assume that $R=(x_{0},y_{0},1)$, where $(x_{0},y_{0})$ satisfies the affine equation (15), in particular $x_{0}\neq 0$. We want compute the divisor $C_{\omega}.T_{R}(C_{\omega})$. Let $f(x,y)=x^{\sqrt{q}}y+\omega xy^{\sqrt{q}}+1$. Since $f_{x}=\omega y^{\sqrt{q}}$ and $f_{y}=x^{\sqrt{q}}$, the tangent line at $R$ is given by
$\omega y_{0}^{\sqrt{q}}x+x_{0}^{\sqrt{q}}y+1=0,$
or, alternatively
$y=-\omega(\frac{y_{0}}{x_{0}})^{\sqrt{q}}x-(\frac{1}{x_{0}})^{\sqrt{q}}$ (20)
because $x_{0}\neq 0$. Take $\sqrt{q}$-th power of (20), then
$y^{\sqrt{q}}=-\omega^{\sqrt{q}}(\frac{y_{0}}{x_{0}})x^{\sqrt{q}}-(\frac{1}{x_{0}}).$ (21)
Substitute (20) and (21) into (15), then
$-\left\{\left(\omega(\frac{y_{0}}{x_{0}})^{\sqrt{q}}+\omega^{\sqrt{q}+1}\frac{y_{0}}{x_{0}}\right)x^{\sqrt{q}+1}+(\frac{1}{x_{0}})^{\sqrt{q}}x^{\sqrt{q}}+\frac{\omega}{x_{0}}x-1\right\}=0.$ (22)
On the other hand, since
$0=\omega\left(\frac{f(x_{0},y_{0})}{x_{0}^{\sqrt{q}+1}}\right)^{\sqrt{q}}=\omega(\frac{y_{0}}{x_{0}})^{\sqrt{q}}+\omega^{\sqrt{q}+1}\frac{y_{0}}{x_{0}}+\frac{\omega}{x_{0}^{\sqrt{q}+1}},$
the coefficient of $x^{\sqrt{q}+1}$ in (22) is $\frac{\omega}{x_{0}^{\sqrt{q}+1}}$. Hence (22) can be written as
$\frac{\omega}{x_{0}^{\sqrt{q}+1}}x^{\sqrt{q}+1}-\frac{1}{x_{0}^{\sqrt{q}}}x^{\sqrt{q}}-\frac{\omega}{x_{0}}x+1$ (23)
$=$ $\frac{\omega}{x_{0}^{\sqrt{q}+1}}\left(x-x_{0}\right)^{\sqrt{q}}\left(x-\frac{x_{0}}{\omega}\right)=0.$ (24)
Since $\omega\neq 1$, $x_{0}\neq\frac{x_{0}}{\omega}$, which means that $R$ is not an inflexion. Hence only $P$ and $Q$ are inflexions of $C_{\omega}$.
Lastly we classify the curves $C_{\omega}$ with $\omega\neq 1$ up to projective equivalence. For $\omega,\omega^{\prime}\in\mathbb{F}_{q}^{\ast}\setminus\operatorname{Ker}(\operatorname{Nm})$, let
$A_{\omega}=\begin{pmatrix}0&1&0\\ \omega&0&0\\ 0&0&1\end{pmatrix}\text{ and }A_{\omega^{\prime}}=\begin{pmatrix}0&1&0\\ \omega^{\prime}&0&0\\ 0&0&1\end{pmatrix}.$
Then two curves $C_{\omega}$ and $C_{\omega^{\prime}}$ are projectively equivalent to each other if and only if there is a matrix $T=(a_{ij})\in GL(3,\mathbb{F}_{q})$ such that
$T^{\ast}A_{\omega}T=\rho A_{\omega^{\prime}}$ (25)
for some $\rho\in\mathbb{F}_{q}^{\ast}$. Since inflexions of both curves are $\{P,Q\}$ that are lie on the line $\{z=0\}$, and
$T_{P}(C_{\omega})\cap T_{Q}(C_{\omega})=T_{P}(C_{\omega^{\prime}})\cap T_{Q}(C_{\omega^{\prime}})=\{(0,0,1)\},$
the line $\{z=0\}$ and the point $(0,0,1)$ are stable by the projective transformation $T$. Hence we may choose $T$ as $\begin{pmatrix}a_{11}&a_{12}&0\\ a_{21}&a_{22}&0\\ 0&0&1\end{pmatrix}.$ Then, comparing $(3,3)$-entries of both sides of (25), we know $\rho=1$. Now we write down (25) componentwise:
$\omega a_{21}^{\sqrt{q}}a_{11}+a_{11}^{\sqrt{q}}a_{21}$ $=0$ (26)
$\omega a_{22}^{\sqrt{q}}a_{12}+a_{12}^{\sqrt{q}}a_{22}$ $=0$ (27)
$\omega a_{21}^{\sqrt{q}}a_{12}+a_{11}^{\sqrt{q}}a_{22}$ $=1$ (28)
$\omega a_{22}^{\sqrt{q}}a_{11}+a_{12}^{\sqrt{q}}a_{21}$ $=\omega^{\prime},$ (29)
where (26)–(29) come from comparison of $(1,1)$,(2,2)22(2,2)( 2 , 2 ), $(1,2)$ and $(2,1)$ components respectively. Take $\sqrt{q}$th power of (26)
$\omega^{\sqrt{q}}a_{21}a_{11}^{\sqrt{q}}+a_{11}a_{21}^{\sqrt{q}}=0,$ (30)
[…]
$(\omega^{\sqrt{q}+1}-1)a_{11}^{\sqrt{q}}a_{21}=0.$
Since $\operatorname{Nm}\omega\neq 1$, we conclude that $a_{11}=0$ or $a_{21}=0$. Similarly, using (27), we have $a_{22}=0$ or $a_{12}=0$. Since $\det T\neq 0$, only possibilities of $T$ are
$T=\begin{pmatrix}a_{11}&0&0\\ 0&a_{22}&0\\ 0&0&1\end{pmatrix}\text{or}\begin{pmatrix}0&a_{12}&0\\ a_{21}&0&0\\ 0&0&1\end{pmatrix}.$
If the first case occurs, then $a_{11}^{\sqrt{q}}a_{22}=1$ by (28) and $\omega a_{22}^{\sqrt{q}}a_{11}=\omega^{\prime}$ by (29), which implies that $\omega=\omega^{\prime}$. If the second case occurs, then
$\omega a_{21}^{\sqrt{q}}a_{12}=1\text{ and }a_{12}^{\sqrt{q}}a_{21}=\omega^{\prime},$
which implies that $\omega\omega^{\prime\sqrt{q}}=1$. ∎
Acknowledgment. The second author was partially supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (2022R1A2C1012291).